# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 3 Randomized arcade processes
An $X$-randomized arcade process (X-RAP), denoted $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ on the partition $\{T_{0},T_{n}\}_{*}$ is a stochastic process of the form
$I_{t}^{(n)}:=D_{t}-\sum_{i=0}^{n}\left(f_{i}(t)D_{T_{i}}-g_{i}(t)X_{i}\right),$ (3.1)
@@
We emphasis that $I_{T_{0}}^{(n)}=X_{0},\ldots,I_{T_{n}}^{(n)}=X_{n}$, so $(I_{t}^{(n)})$ is a strong stochastic interpolator between the random variables $X_{0},\ldots,X_{n}$. The signal function $(S_{t}^{(n)})$ is independent of the noise process $(A_{t}^{(n)})$ since $X$ is assumed independent of $(D_{t})_{t\in[T_{0},T_{n}]}$.
@@
A related but distinct class of processes, introduced in [22] and called random $n$-bridges, shares some characteristics with RAPs. These processes, defined weakly, match given probability measures instead of random variables. They are constructed in a similar fashion as stochastic bridges, that is, by conditioning a stochastic process to match given distributions at given times. In special cases, the law of a RAP satisfies the conditions for the RAP to be a random $n$-bridge. For instance, the RAP obtained by randomizing the stitched Brownian bridge, using the same interpolating coefficients for the signal function as the ones used in the noise process, has a law that satisfies the conditions for the RAP to be a random $n$-bridge. However, any other RAP driven by Brownian motion is not a random $n$-bridge. Conversely, certain random $n$-bridges cannot have the same law as a RAP. In this paper, we chose not to consider the case in which the driver underlying a RAP jumps. But if we did, just for the sake of comparison, a random $n$-bridge built using a gamma process will not match the law of any RAP driven by a gamma process, since APs are sums, not products. The development of alternative arcade processes driven by (pure) jump processes will be pursued in future research. It requires a different construction that does not rely on the covariance function, similar to the change needed to go from Gaussian bridges to Lévy bridges [12] or Markov [20] bridges.
@@
Let $X=(X_{0},\ldots,X_{5})$ be a vector of independent and uniform $\mathcal{U}\,(\{-1,1\})$ random variables, and $Y=(Y_{0},\ldots,Y_{5})$ be another vector of random variables such that $Y_{0}\sim\mathcal{U}\,(\{-1,1\})$, $Y_{i}=-Y_{i-1}$ for $i=1,\ldots,5$. Let $(A_{t}^{(5)})$ be an AP with elliptic interpolation coefficients driven by Brownian motion, $g_{i}=f_{i}$ for $i=0,\ldots,5$, and $(I_{t}^{(5)})$, $(\tilde{I}_{t}^{(5)})$ its associated $X$-RAP and $Y$-RAP, respectively. Although the same driver and interpolating coefficients for both RAPs are the same, and the vectors $X$ and $Y$ have the same marginal distributions, the paths of $(I_{t}^{(5)})$ and $(\tilde{I}_{t}^{(5)})$ are different.
Figure 6: Paths simulation of $(I_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 3.3.
Figure 7: Paths simulation of $(\tilde{I}_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 3.3.
Besides its main purpose of interpolating in the strong sense, a RAP can also be used to mimic another stochastic process. Let $(Y_{t})_{t\in[T_{0},T_{n}]}$ be a sample-continuous stochastic process. For instance, if $\{T_{0},T_{n}\}_{*}$ is the equidistant partition of an interval $[a,b]$, $X=(Y_{T_{0}},\ldots,Y_{T_{n}})$, and $\{f_{0},\ldots,f_{n}\}=\{g_{0},\ldots,g_{n}\}$ are the piecewise linear interpolating coefficients, then for almost all $\omega\in\Omega$, $\sup\limits_{t\in[a,b]}A_{t}^{(n)}(\omega)\rightarrow 0$, and $\sup\limits_{t\in[a,b]}S_{t}^{(n)}(\omega)\rightarrow Y_{t}(\omega)$ as $n\rightarrow\infty$. Hence, in that case,
[…]
@@
Example 3.4.
Let $[a,b]=[0,10]$, and $\{T_{0},T_{n}\}_{*}$ be its equidistant partition. If $(Y_{t})$ is a fractional Brownian motion (fBm) and $X=(Y_{T_{0}},\ldots,Y_{T_{n}})$, then the paths of an $X$-randomized stitched Brownian arcade are similar to the one of $(Y_{t})$ when $n$ is large enough.
Figure 8: Path of the fBm $(Y_{t})$ , c.f., Example 3.4.
Figure 9: Path of an $X$-RAP mimicking the fBm $(Y_{t})$, c.f., Example 3.4.
@@
$\mu_{I}(t):=\mathbb{E}[I_{t}^{(n)}]=\mu_{S}(t)+\mu_{A}(t),$ (3.5)
$\sigma_{I}^{2}(t):=\mathrm{Var}{[I_{t}^{(n)}]}=\sigma^{2}_{S}(t)+\sigma^{2}_{A}(t),$ (3.6)
$K_{I}(s,t):=\mathrm{Cov}[{I_{s}^{(n)},I_{t}^{(n)}}]=K_{S}(s,t)+K_{A}(s,t)=\sum\limits_{i=0}^{n}\sum\limits_{j=0}^{n}g_{i}(t)g_{j}(s)\mathrm{Cov}[X_{i},X_{j}]+K_{A}(s,t).$ (3.7)
The proof follows from the proof of Prop. 2.7. We introduce terminologies similar to the ones from the previous section.
@@
Aside of simple cases, an $X$-RAP is not Markovian since no constraints are applied to the joint distribution of $X$. Hence, we introduce a relaxed notion of Markov property better suited to RAPs.
@@
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval, and $\tau_{0}<\tau_{1}<\ldots<\infty$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite, or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called $\tau$-conditionally-Markov if
@@
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, and $\tau_{m(s)}=\max\limits_{i\in\mathbb{N}}\{\tau_{i}\,|\,\tau_{i}\leqslant s\}$.
Now, we give a result on the conditional-Markov property of Gaussian RAPs similar to the one on the Markov property of Gaussian APs.
@@
Let $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}=(S_{t}^{(n)}+A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ be a Gaussian $X$-RAP on $\{T_{0},T_{n}\}_{*}$. Then $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov if the following conditions are all satisfied:
@@
    The AP $(A_{t}^{(n)})$ is Markov, i.e., $K_{A}(s,t)=\sum\limits_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t)$.
@@
Let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov if and only if
@@
where $m(s_{k}):=\max\limits\{i\in\mathbb{N}\,|\,T_{i}\leqslant s_{k}\}$. In the following, we will refer to $m(s_{k})$ by $m$ since $s_{k}$ is fixed.
@@
by Eq. (3.9). We also know that $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(X_{0},\ldots,X_{m})$ by the definition of the $X$-RAP, and $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(I_{s_{j+1}}^{(n)},\ldots,I_{s_{k}}^{(n)},I_{t}^{(n)})$ since $(A_{t}^{(n)})$ is Markov. We conclude that
@@
which means we can assume that $s_{1},\ldots,s_{k}$ are all in the same sub-interval $(T_{m},T_{m+1})$.
Let us define $a_{m}(\cdot):=\sum\limits_{i=0}^{m}g_{i}(\cdot)X_{i}$, and
$\Delta_{q}:=\sum_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum_{i=1}^{k}c_{i,q}g_{m+1}(s_{i})X_{m+1}+\sum_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)},\quad q=1,\ldots,k-1,$ (3.14)
where the coefficients $(c_{i,q})$ are chosen such that
@@
This guarantees the following, where the notation ”$\,|\,(X_{0},\ldots,X_{m})$” means conditional on $(X_{0},\ldots,X_{m})$).
@@
    $\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right]=\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},\Delta_{1},\ldots,\Delta_{k-1},I_{s_{k}}^{(n)}\right]$.
@@
    $(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$ is a Gaussian vector. To see this, we observe that for all $q=1,\ldots k-1$,
@@
    where we used Eq. (3.10). Hence, $\Delta_{q}=\sum\limits_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum\limits_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum\limits_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)}$ for all $q=1,\ldots,k-1$, which implies that $(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$ is a Gaussian vector.
@@
    $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$, since $\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0$.
@@
    $A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$, since $(A_{t}^{(n)})$ is Markov.
@@
$I_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})\text{ and }I_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m}).$ (3.17)
Since $(I_{t}^{(n)})=(S_{t}^{(n)}+A_{t}^{(n)})$, and $(S_{t}^{(n)})\perp\!\!\!\perp(A_{t}^{(n)})$, we have 3.17 if
$A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m}),$ (3.18)
which is guaranteed by conditions 3.15. ∎
@@
This makes standard RAPs automatically conditionally-Markov.
@@
It is important to note that the conditional-Markov property is not symmetric in time. Define $\mathcal{G}_{t}^{I}=\sigma(I_{u}^{(n)}\,|\,t\leqslant u\leqslant T_{n})$, and let $s<t$ in $[T_{0},T_{n}]$. Then, to obtain
@@
where $k(t)=\min\limits\{i\in\mathbb{N}\,|\,T_{i}\geqslant t\}$, Condition (3.9) needs to be replaced by
@@
We give an example of a non-standard $X$-RAP on $[T_{0},T_{2}]$ that is $\{T_{0},T_{2}\}_{*}$-conditionally-Markov, where $X_{0}\overset{\mathcal{L}}{=}X_{1}\overset{\mathcal{L}}{=}X_{2}\overset{\mathcal{L}}{=}\textrm{Uni}\,(\{-1,1\})$ are mutually independent. Consider the interpolating coefficients
@@
Outside of the considered intervals, the functions $g_{i}$ may take any values as long as they remain interpolating coefficients. Theorem 3.8 does not impose a condition on $g_{0}$. For example, we could choose $g_{i}=f_{i}$ outside the above intervals. Hence, all three conditions are fulfilled and this $X$-RAP is $\{T_{0},T_{2}\}_{*}$-conditionally-Markov. As we can see from the paths-simulation below, this process is different from a randomized stitched Brownian arcade on the second arc. The noise has been diminished to make the paths more informative. Simulating the signal function highlights the following: $X_{0}$ will determine the fate of the signal function on $[T_{1},T_{2}]$ since this RAP remembers the previously matched random variables when changing arc. On the first arc, where the process is simply a randomized Brownian bridge, to go from $X_{0}=-1$ to $X_{1}=-1$, for instance, there is only one way—a straight line. On the second arc, to go from $X_{1}=-1$ to $X_{2}=-1$, there are two ways. The evolution of the signal function is determined by the value that $X_{0}$ takes. This is illustrated by the paths of the signal function below: the blue path and the green path both take value $-1$ at $T_{1}$ and value $1$ at $T_{2}$, but have different values in $T_{0}$. Hence they differ on $[T_{1},T_{2}]$, as observed.
Figure 10: Paths simulation of the signal function of a non-standard $X$-RAP, where $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$ , c.f., Example 3.12.
Figure 11: Paths simulation of a non-standard $X$-RAP, where the noise is rescaled by 0.3 and $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$, c.f., Example 3.12.

### v3

@@ 3 Randomized arcade processes
An $X$-randomised arcade process ($X$-RAP) on the partition $\{T_{0},T_{n}\}_{*}$, denoted $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$, is a stochastic process of the form
$I_{t}^{(n)}\mathrel{\mathop{\ordinarycolon}}=D_{t}-\sum_{i=0}^{n}\left(f_{i}(t)D_{T_{i}}-g_{i}(t)X_{i}\right),$ (3.1)
@@
We emphasise that $I_{T_{0}}^{(n)}=X_{0},\ldots,I_{T_{n}}^{(n)}=X_{n}$, so $(I_{t}^{(n)})$ is a strong stochastic interpolator between the random variables $X_{0},\ldots,X_{n}$. The signal function $(S_{t}^{(n)})$ is independent of the noise process $(A_{t}^{(n)})$ since $X$ is assumed to be independent of $(D_{t})_{t\in[T_{0},T_{n}]}$.
@@
A related but distinct class of processes, introduced in [24] and called random $n$-bridges, shares some characteristics with RAPs. These processes, defined weakly, match given probability measures instead of random variables. They are constructed in a similar fashion as stochastic bridges, that is, by conditioning a stochastic process to match given distributions at given times. In special cases, the law of a RAP satisfies the conditions for the RAP to be a random $n$-bridge. For instance, the RAP obtained by randomising the stitched Brownian bridge, using the same interpolating coefficients for the signal function as the ones used in the noise process, has a law that satisfies the conditions for the RAP to be a random $n$-bridge. However, any other RAP driven by Brownian motion is not a random $n$-bridge. Conversely, certain random $n$-bridges cannot have the same law as a RAP. In this work, we chose not to consider the case in which the driver underlying a RAP jumps. But if we did, just for the sake of comparison, a random $n$-bridge built using a gamma process would not match the law of any RAP driven by a gamma process, since APs are sums, not products. The development of alternative arcade processes driven by (pure) jump processes will be pursued in future research. It requires a different construction that does not rely on the covariance function, similar to the change needed to go from Gaussian bridges to Lévy bridges [13] or Markov bridges [22] .
@@
Let $X=(X_{0},\ldots,X_{5})$ be a vector of independent and uniform $\mathcal{U}\,(\{-1,1\})$ random variables and $Y=(Y_{0},\ldots,Y_{5})$ be another vector of random variables such that $Y_{0}\sim\mathcal{U}\,(\{-1,1\})$, $Y_{i}=-Y_{i-1}$ for $i=1,\ldots,5$. Let $(A_{t}^{(5)})$ be an AP with elliptic interpolation coefficients driven by Brownian motion, $g_{i}=f_{i}$ for $i=0,\ldots,5$, and $(I_{t}^{(5)})$, $(\tilde{I}_{t}^{(5)})$ its associated $X$-RAP and $Y$-RAP, respectively. Although the same driver and interpolating coefficients for both RAPs are the same, and the vectors $X$ and $Y$ have the same marginal distributions, the paths of $(I_{t}^{(5)})$ and $(\tilde{I}_{t}^{(5)})$ are different.
Figure 1: Paths simulation of $(I_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\mid i=0,1,\ldots,5\}$, cf., Example 3.18.
Figure 2: Paths simulation of $(\tilde{I}_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\mid i=0,1,\ldots,5\}$, cf., Example 3.18.
As a side remark, we note here that besides its main purpose of interpolating in the strong sense, a RAP can also be used to mimic a stochastic process. Let $(Y_{t})_{t\in[T_{0},T_{n}]}$ be a sample-continuous stochastic process. For instance, if $\{T_{0},T_{n}\}_{*}$ is the equidistant partition of an interval $[a,b]$, $X=(Y_{T_{0}},\ldots,Y_{T_{n}})$, and $\{f_{0},\ldots,f_{n}\}=\{g_{0},\ldots,g_{n}\}$ are the piecewise linear interpolating coefficients, then for almost all $\omega\in\Omega$, $\sup\limits_{t\in[a,b]}A_{t}^{(n)}(\omega)\rightarrow 0$, and $\sup\limits_{t\in[a,b]}\mathinner{\!\left\lvert S_{t}^{(n)}(\omega)-Y_{t}(\omega)\right\rvert}\rightarrow 0$ as $n\rightarrow\infty$. Hence, in this case,
$\sup\limits_{t\in[a,b]}\mathinner{\!\left\lvert I_{t}^{(n)}-Y_{t}\right\rvert}=\sup\limits_{t\in[a,b]}\mathinner{\!\left\lvert S_{t}^{(n)}+A_{t}^{(n)}-Y_{t}\right\rvert}\rightarrow 0,$ (3.4)
@@
$\mu_{I}(t)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[I_{t}^{(n)}]=\mu_{S}(t)+\mu_{A}(t),$ (3.5)
$\sigma_{I}^{2}(t)\mathrel{\mathop{\ordinarycolon}}=\mathrm{Var}{[I_{t}^{(n)}]}=\sigma^{2}_{S}(t)+\sigma^{2}_{A}(t),$ (3.6)
$K_{I}(s,t)\mathrel{\mathop{\ordinarycolon}}=\mathrm{Cov}[{I_{s}^{(n)},I_{t}^{(n)}}]$
$\hskip 42.67912pt=K_{S}(s,t)+K_{A}(s,t)=\sum\limits_{i=0}^{n}\sum\limits_{j=0}^{n}g_{i}(t)g_{j}(s)\mathrm{Cov}[X_{i},X_{j}]+K_{A}(s,t).$ (3.7)
The proof follows from the proof of Prop. 2.4. We introduce terminologies similar to those in the previous section.
@@
4 Conditionally Markov RAPs
Aside of simple cases, an $X$-RAP is not Markovian since no constraints are applied to the joint distribution of $X$. Hence, we introduce a relaxed notion of the Markov property that is better suited to RAPs.
@@ 4 Conditionally Markov RAPs
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval and $\tau_{0}<\tau_{1}<\ldots<\infty$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called $\tau$-conditionally Markov if
@@
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, and $\tau_{m(s)}=\max\limits_{i\in\mathbb{N}}\{\tau_{i}\mid\tau_{i}\leqslant s\}$.
Now, we give a result on the conditionally-Markov property of Gaussian RAPs similar to the one on the Markov property of Gaussian APs.
@@
Let $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}=(S_{t}^{(n)}+A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ be a Gaussian $X$-RAP on $\{T_{0},T_{n}\}_{*}$. Then $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally Markov if the following conditions are all satisfied:
@@
    The AP $(A_{t}^{(n)})$ is Markov, i.e., $K_{A}(s,t)\\ =\sum\limits_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t)$.
@@
Consider Theorem 1.12 in [20], and let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally Markov if and only if
@@
where $m(s_{k})\mathrel{\mathop{\ordinarycolon}}=\max\limits\{i\in\mathbb{N}\mid T_{i}\leqslant s_{k}\}$. In the following, we will refer to $m(s_{k})$ by $m$ since $s_{k}$ is fixed.
@@
by Eq. (4.2). We also know that $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(X_{0},\ldots,X_{m})$ by the definition of the $X$-RAP, and therefore, in particular, $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(S_{s_{j+1}}^{(n)},\ldots,S_{s_{k}}^{(n)},S_{t}^{(n)})$. Furthermore, since $(A_{t}^{(n)})$ is Markov and $A_{T_{m}}^{(n)}=0$, we have $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(A_{s_{j+1}}^{(n)},\ldots,A_{s_{k}}^{(n)},A_{t}^{(n)})$, and hence $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(I_{s_{j+1}}^{(n)},\ldots,I_{s_{k}}^{(n)},I_{t}^{(n)})=(S_{s_{j+1}}^{(n)}+A_{s_{j+1}}^{(n)},\ldots,S_{s_{k}}^{(n)}+A_{s_{k}}^{(n)},S_{t}^{(n)}+A_{t}^{(n)})$. We conclude that
@@
which means, we can assume that $s_{1},\ldots,s_{k}$ are all in the same sub-interval $(T_{m},T_{m+1})$.
Let us define $a_{m}(\cdot)\mathrel{\mathop{\ordinarycolon}}=\sum\limits_{i=0}^{m}g_{i}(\cdot)X_{i}$, and
$\Delta_{q}\mathrel{\mathop{\ordinarycolon}}=\sum_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum_{i=1}^{k}c_{i,q}g_{m+1}(s_{i})X_{m+1}+\sum_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)},$ (4.7)
for $q=1,\ldots,k-1$, where the coefficients $(c_{i,q})$ are chosen such that
@@
This guarantees the following, where the notation $\mid(X_{0},\ldots,X_{m})$" means conditional on $(X_{0},\ldots,X_{m})$.
@@
    $\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right]\\ =\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},\Delta_{1},\ldots,\Delta_{k-1},I_{s_{k}}^{(n)}\right]$.
@@
    $(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})$ is a Gaussian vector. To see this, we observe that for all $q=1,\ldots k-1$,
@@
    where we used Eq. (4.3). Hence, $\Delta_{q}=\sum\limits_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum\limits_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum\limits_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)}$ for all $q=1,\ldots,k-1$, which implies that $(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})$ is a Gaussian vector.
@@
    $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})$, since $\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0$.
@@
    $A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})$, since $(A_{t}^{(n)})$ is Markov.
@@
$I_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})\text{ and }I_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m}).$ (4.10)
Since $(I_{t}^{(n)})=(S_{t}^{(n)}+A_{t}^{(n)})$, and $(S_{t}^{(n)})\perp\!\!\!\perp(A_{t}^{(n)})$, we have (4.10) if
$A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\mid(X_{0},\ldots,X_{m}),$ (4.11)
which is guaranteed by conditions (4.8). ∎
@@
This makes standard RAPs automatically conditionally Markov.
@@
It is important to note that the conditionally-Markov property is not symmetric in time. Define $\mathcal{G}_{t}^{I}=\sigma(I_{u}^{(n)}\mid t\leqslant u\leqslant T_{n})$, and let $s<t$ in $T_{0},T_{n}]$. Then, to obtain
@@
where $k(t)=\min\limits\{i\in\mathbb{N}\mid T_{i}\geqslant t\}$, Condition (4.2) needs to be replaced by
@@
We give an example of a non-standard $X$-RAP on $T_{0},T_{2}]$ that is $\{T_{0},T_{2}\}_{*}$-conditionally Markov, where $X_{0}\overset{\mathcal{L}}{=}X_{1}\overset{\mathcal{L}}{=}X_{2}\overset{\mathcal{L}}{=}\textrm{Uni}\,(\{-1,1\})$ are mutually independent. Consider the interpolating coefficients
@@
Outside of the considered intervals, the functions $g_{i}$ may take any values as long as they remain interpolating coefficients. Theorem 4.22 does not impose a condition on $g_{0}$. For example, we could choose $g_{i}=f_{i}$ outside the above intervals. Hence, all three conditions are fulfilled and this $X$-RAP is $\{T_{0},T_{2}\}_{*}$-conditionally Markov. As we can see from the paths-simulation below, this process is different from a randomised stitched Brownian arcade on the second arc. The noise has been diminished to make the paths more informative. Simulating the signal function highlights the following: $X_{0}$ will determine the fate of the signal function on $T_{1},T_{2}]$, since this RAP remembers the previously matched random variables when changing arcs. On the first arc, where the process is simply a randomised Brownian bridge, to go from $X_{0}=-1$ to $X_{1}=-1$, for instance, there is only one way—a straight line. On the second arc, to go from $X_{1}=-1$ to $X_{2}=-1$, there are two ways. The evolution of the signal function is determined by the value that $X_{0}$ takes. This is illustrated by the paths of the signal function below: the blue path and the green path both take value 1-1 at $T_{1}$ and value $1$ at $T_{2}$, but have different values in $T_{0}$. Hence, they differ on $T_{1},T_{2}]$, as observed.
Figure 3: Paths simulation of the signal function of a non-standard $X$-RAP, where $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$, cf., Example 4.26.
Figure 4: Paths simulation of a non-standard $X$-RAP, where the noise is rescaled by 0.3 and $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$, cf., Example 4.26.